TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.MircoCoccoli on 2006-12-20 - 12:01

RAT Preparation Cool Down sect 78 - Dec 20th

PDF RAT_20061220.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x11x10/RAT_20061220.pdf

Next RAT meeting Tuesday, January 9th 8:30 @ P8 2889-R-009

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 15.12.06 DFBXG non-conformity on the current lead heater cables (P.Chambouvet-J-C.Guillaume)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
13 kA EE Systems
  • 16.12.06 Calibration of the Arc Detectors on the 13 kA EE System (K.Dahlerup-Petersen)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

MircoCoccoli - 20 Dec 2006 - 11:11

Edit, attach file or add comment


Created by Main.IvanMoyaMartinez on 2006-12-20 - 10:57

The Flags are Up

Polarity Check done in RR57-UA56

JPEG P1000172.JPG:
P1000172.JPG

JPEG P1000173.JPG:
P1000173.JPG

JPEG P1000174.JPG:
P1000174.JPG

JPEG P1000157.JPG:
P1000157.JPG

JPEG P1000154.JPG:
P1000154.JPG

JPEG P1000155.JPG:
P1000155.JPG

Comments:

  • Updated. Cables labeled following the ECR(EDMS 804188) -- Main.IvanMoyaMartinez - 09 Jan 2007 - 11:00

Edit, attach file or add comment


Created by Main.RobertoSaban on 2006-12-20 - 09:52

Incident Summary Report - Short on the MB circuit during TP4B

During the Electrical Quality Assurance (TP4B) of sector 78 a short to ground was identified. The short chronological report of the events is attached.

PDF short_21R7.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/short_21R7.pdf

PowerPoint FSavary.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/FSavary.ppt
Report to MARIC of the repair sequence
PowerPoint DBozzini.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/DBozzini.ppt
Report to MARIC of the ELQA tests

Comments:

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-12-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login